This is all one problem. I'm having trouble with this: In 1997, the population of the United States was about 266 million and the exponential growth rate 0.9% per year.
a) find the exponential growth function
b) what will the population be in 2002?
d) when will the population be doubled?

Answer by stanbon(48516)

This is all one problem. I'm having trouble with this: In 1997, the population of the United States was about 266 million and the exponential growth rate 0.9% per year.
a) find the exponential growth function
pop=266(1.09)^x million where x is the number of years after 1997)
-------------------------------------------
b) what will the population be in 2002?
P(5)=266(1.09)^5 = 409.27 million
-------------------------------------------
d) when will the population be doubled?
2*266=266(1.09)^x
2 = 1.09^x
x= log2/log1.09
x = 8.04 years
---------------
Year = 1997+8.04 = 2005.04
So, the answer is either 2005 or 2006 depending on what you do with
that 0.04 years.
===================
Cheers,
Stan H.